OpenStudy (anonymous):

@misty1212 Are you familiar with Intro Topology Proofs?

OpenStudy (misty1212):

HI NO

OpenStudy (misty1212):

although i have seen some you got one in mind?

OpenStudy (anonymous):

:< yeah a proof on my HW i have to do.

OpenStudy (misty1212):

post it i understand anything i will say so

OpenStudy (misty1212):

*IF i understand

OpenStudy (anonymous):

Let f:R->R be defined by: f(x) = 1/x for x>0 and 0 for x=< 0 Prove that the graph G_f is a closed subset of (R^2,d), but that f is not continous

OpenStudy (anonymous):

g_f is defined to be the set {(x,y) | (x,y) is in (R x R) where y = f(x)}

OpenStudy (anonymous):

@iGreen. Could you help?